# A statistical test for network similarity

### v1

@@ 4.2.2 Real-world graphs
We end our computational tests with real-world graphs of varying sizes. We conducted our tests on the following (directed and undirected) graphs:
  • email-Eu-core network [15, 37] (directed, $N=1,005$)
  • Wikipedia vote network [14, 13] (directed, largest connected component only $N=7,066$)
@@
Table 7: Nodes removal experiment for five real-world graphs
Percent Removed 0 0.5 1 5 10 25
email-Eu-Core (Directed) 0 (1) 1e-4 (1) 0.002 (0.495) 0.010 (0.996) 0.011 (1e-23) 0.044 (1e-297)
Wikipedia (Directed) 0 (1) 0.00016 (0.896) 0.00023 (0.484) 0.0018 (4e-32) 0.0067 (0) 0.0116 (0)
W & S power grid 0 (1) 1e-5 (1) 1e-6 (1) 1e-5 (1) 1e-4 (1) 1e-4 (0.453)
Karate club 0 (1) 0 (1) 0 (1) 0.026 (0.989) 0.026 (0.993) 0.123 (0)
Facebook egonets 0 (1) 1e-4 (0.684) 1e-4 (0.540) 0.003 (1e-27) 0.007 (1e-149) 0.027 (0)
Table 8: Edges removal experiment for five real-world graphs
Percent Removed 0 0.5 1 5 10 25
email-Eu-Core (Directed) 0 (1) 0.001 (0.773) 0.004 (0.001) 0.016 (1e-56) 0.033 (1e-230) 0.089 (0)
Wikipedia (Directed) 0 (1) 0.00047 (1) 0.0009 (3e-9) 0.0048 (4e-252) 0.0099 (0) 0.0256 (0)
W & S power grid 0 (1) 1e-5 (1) 1e-5 (1) 1e-4 (1) 1e-4 (0.839) 1e-4 (0.027)
Karate club 0 (1) 0 (1) 0.010 (1) 0.021 (1) 0.087 (0.028) 0.226 (1e-13)
Facebook egonets 0 (1) 0.001 (0.005) 0.003 (1e-31) 0.012 (0) 0.023 (0) 0.057 (0)
5 Conclusion
@@
While our metric is unarguably accurate, we observe that in many cases the distance between networks may be small but significant. For this reason, we suggest that our metric only be used in conjunction with domain expertise (e.g., knowledge of what is a typical graph variation), when performing anomaly or change-point detection tasks.
Future work should focus on the validity of the p-values. Indeed, while we notice a clear monotonic decrease in p-values in step with increases in network dissimilarity, we would like to formally examine their statistical validity.

### v2

@@ 4.2.2 Real-world graphs
We also conduct computational tests on real-world graphs of varying sizes. We conducted our tests on the following (directed and undirected) graphs:
  • email-Eu-core network [15, 38] (directed, largest connected component only, $N=986$)
  • Wikipedia vote network [14, 13] (directed, largest connected component only, $N=7,066$)
@@
Table 11: Nodes removal experiment for real-world graphs
Percent Removed 0 1 5 10 25
email-Eu-Core (Directed) 0 0.002 na na na
Wikipedia (Directed) 0 na na na na
W & S power grid 0 na na na na
Facebook egonets 0 0.000 0.003 0.007 na
Table 12: Edges removal experiment for real-world graphs
Percent Removed 0 1 5 10 25
email-Eu-Core (Directed) 0 na na na na
Wikipedia (Directed) 0 na na na na
W & S power grid 0 na na na na
Facebook egonets 0 0.003 na na na
Here too, the real world graphs are extremely sparse graph. The removal of a very small number of nodes or edges does not affect this reality. Therefore, the graph structure is largely unaffected. Our test results confirm this reality. Once again, to demonstrate how the node/edge removal in our experiments did not modify network structures, beyond causing disconnections, we show the evolution of each graph’s density. In Table 13 and Table 14 we clearly observe that the node/edges removal have only minor effects on network structure, beyond disconnection (na).
Table 13: Densities under node removal experiment for real-world graphs
Percent Removed 0 1 5 10 25
email-Eu-Core (Directed) 0.026 0.027 na na na
Wikipedia (Directed) 0.002 na na na na
W & S power grid 0.001 na na na na
Facebook egonets 0.011 0.011 0.011 na na
Table 14: Densities under edge removal experiment for real-world graphs
Percent Removed 0 1 5 10 25
email-Eu-Core (Directed) 0.026 na na na na
Wikipedia (Directed) 0.002 na na na na
W & S power grid 0.001 na na na na
Facebook egonets 0.011 0.011 na na na
4.2.3 Node and vertex additions
In order to examine our test results on the network structure modification of very sparse graphs, we add nodes and random connections to other nodes on the graph and perform our test. For the experiments presented in Table 15, we revisit the two synthetic (CM) and the four real-world graphs. These graphs were too sparse to be studied using node or edge removal, since they quickly, sometimes instantly, became disconnected. Instead of removing edges or nodes, we add a pre-determined number of nodes to each graph. This number is a percentage of the number of pre-existing nodes. We also add the same percentage of connections to other randomly selected nodes in the graph.
Table 15: Node and edge addition
Percent (of nodes and connections) Added 0 1 5 10 25
CM 3400 (Directed) 0 0.008 0.594 0.997 0.998
email-Eu-core (Directed) 0 0.003 0.14 0.486 0.88
Wiki-Vote (Directed) 0 0.014 0.752 0.916 0.986
CM 3400 0 0.004 0.374 0.972 0.999
W & S power grid 0 0.006 0.483 0.993 0.999
Facebook egonets 0 0.006 0.363 0.811 0.956
Finally, to contrast these experiments to node or edge removal ones, we show the evolution of graph densities in Table 16. These final experiments highlight how our test will capture even a very small change in network structure. Along with all other experiments on the very sparse CM and real-world graphs, they also highlight how the significance of a change in structure is relative to the graph’s initial state, not a universal scale, like a fixed percent change in the number of nodes or edges.
Table 16: Densities under node and edge addition experiments
Percent (of nodes and connections) Added 0 1 5 10 25
CM 3400 (Directed) 0.001 0.001 0.005 0.019 0.098
email-Eu-core (Directed) 0.026 0.026 0.029 0.04 0.114
Wiki-Vote (Directed) 0.002 0.002 0.007 0.02 0.099
CM 3400 0.001 0.001 0.005 0.019 0.098
Watts-Strogatz 0.001 0.001 0.005 0.019 0.098
Fbook 0.011 0.011 0.015 0.027 0.104
5 Discussion and conclusions
@@
While our metric is unarguably accurate, we observe that in many cases the distance between networks may be small. For this reason, we suggest that our metric only be used in conjunction with domain expertise (e.g., knowledge of what is a typical graph variation) and a thorough examination of the graph’s structure, when performing anomaly or change-point detection tasks. We believe that the significance of a distance is more a domain than a statistical question. In particular, our experiments with the very sparse (CM and real-world) graphs demonstrate that the significance in a change is relative to each graph’s initial state, not an arbitrary variation in the number of edges or connections.
Future work should focus on change-point and anomaly detection and on the validity of the $p$-values. Indeed, while we notice a clear monotonic decrease in $p$-values in step with increases in network dissimilarity, we would like to formally examine their statistical validity.